A rectangular parking lot has an area of 2400ft^2. If the width is 20 feet less than the length, what are the dimensions of the parking lot?
Answer by checkley79(3341)

A=LW
W=L-20
2,400=L(L-20)
2,400=L^2-20L
L^2-20L-2,400=0
(L-60)(L+40)=0
L-60=0
L=60 ANS. FOR THE LENGTH.
W=60-20=40 ANS. FOR THE WIDTH.
PROOF:
2,400=60*40
2,400=2,400